I am a little confused on these two problems from the book. We are working on squared roots and am confused. The problems are story problems.
The radius of a sphere is given by the formula r = (0.75v over pie)^1/3 Where V is its volume. Find the radius of a spherical tank that has a volume at 32pie over 3 cubic meters.
An investment of $10,000 in 1997 in the Spartan Investment Grade bond Fund grew to $14,309.61 in 2002. Use the formula r = (s over p)^1/n - 1 to find the 5 year average annual return.
Can someone please help?

Answer by stanbon(75887)

The radius of a sphere is given by the formula r = (0.75v over pie)^1/3 Where V is its volume. Find the radius of a spherical tank that has a volume at 32pie over 3 cubic meters.
r = [(0.75(32(pi)/3/(pi)]^(1/3)
Cancel the pi's
r = [(3/4)(32)/3]^(1/3)
r = 8^(1/3)
r = 2 meters
----------------
An investment of $10,000 in 1997 in the Spartan Investment Grade bond Fund grew to $14,309.61 in 2002. Use the formula r = (s over p)^1/n - 1 to find the 5 year average annual return.
r = (14,309.61/10,000)^(1/n)-1
r = 1.43061^(1/5)-1
r = 1.0742-1
r = 0.0742...
================
Cheers,
Stan H.